Question. points A force F-200 N acts to pull a fish box of mass 40 kg on a rough horizontal surface of up-0.25. What is the acceleration of


the box (in m/s²)? (g=10 m/s²)


1
Expert's answer
2021-05-23T16:34:40-0400

Let's apply Newton's Second Law of Motion:


"F_{pull}-F_{fr}=ma,""F_{pull}-\mu mg=ma,""a=\dfrac{F_{pull}-\mu mg}{m},""a=\dfrac{200\ N-0.25\cdot40\ kg\cdot10\ \dfrac{m}{s^2}}{40\ kg}=2.5\ \dfrac{m}{s^2}."